1. On October 20, Cullumber sells merchandise with a cost of $3,840 to Blossom for $6,400 cash. On October 25, Cullumber received returned goods from Blossom and refunds the amount in cash for items that had a selling price of $640 and a cost of $384.
Use a tabular summary to record the transactions for Cullumber Company using a perpetual inventory system. Include margin explanations for the changes in revenues and expenses. (Enter negative amounts using either a negative sign preceding the number e.g. -45 or parentheses e.g. (45).)

2. Pharoah Company provides this information for the month ended October 31, 2022: sales on credit $336,000, cash sales $168,000, sales discounts $5,600, and sales returns and allowances $21,280.
Prepare the sales section of the income statement based on this information.

3. Pronghorn Corp reported net sales of $256,000, cost of goods sold of $148,480, operating expenses of $56,000, net income of $28,160, beginning total assets of $511,200, and ending total assets of $639,200.

Calculate profit margin and gross profit rate.

Profit margin enter profit margin in percentages %
Gross profit rate enter gross profit rate in percentages %
Answer #1

1.) Cucumber company: -

Sales for cash (A) $6400

Sales Return for cash (B) $640

Net Sales (A-B) $5760

Cost of Goods Sold (C) $3840

Goods Returned (D) $384

Net COGS (C-D) $3456  

Journals: -

Cash A/c Dr. $5760

To Sales A/c $5760

(Being Net sales recorded in books)

COGS A/c Dr. $3456

To Cash A/c $3456

(Being goods sold for cash)

Gross Profit / Profit Margin $(5760-3456) = $2304

Gross Profit Rate / Profit Margin = 2304/5760*100 = 40%

Note: - sales return have been reduced from total sales, since net sales have to be recorded in books and consider while calculating gross or profit margins. Similarly, return goods cost have been reduced to get net COGS.

2.) Pharoah Company;' -

Sales Section of the Income Statement: -

Sales on Credit (i) $336000

Cash Sales (ii) $168000

Total Sales (i+ii) = (A) $504000

Sales Discount (B) $5600

Sales Retuns & Allowances (C) $21280

Net Sales (A-B-C) $477120

3.) Pronghorn Corp: -

Profit Margin and Gross Profit rate: -

Net Sales (A) $256000

Cost of Goods Sold (B) $148480

Gross Profit (A-B) = $107520

Gross Profit rate = 107520/256000*100 = 42%

Gross Profit (A)    $107520

Operating Expenses (B)   $56000

Other Expenses (A-B-C) $23360

Net Income/Profit Margin (C) $28160

Profit Margin Rate = 28160/256000*100 = 11%

Note : -Since Net Income is $28160, there has to be a other expenses which have been identified of $23360